# Enhancing Real-World Complex Network Representations with Hyperedge Augmentation

### v1

@@ Appendix A Full Statistics of Evaluation Datasets
Table 7. Statistics of all 26 datasets built for HyperAug evaluation.
@@ Appendix B Dataset and Task Specifications
B.1. MUSAE Datasets
B.1.1. MUSAE-GitHub
@@
This is a large social network of GitHub developers introduced in MUSAE (Rozemberczki et al., 2021). Nodes represent developers on GitHub, and edges are mutual follower relationships. HyperAug constructs the virtual hyperedges as mutually following developer groups that contain at least 3 developers (i.e., maximal cliques with sizes of at least 3). We enable the option to use either the 4,005-dimensional raw node features extracted based on the location, repositories starred, employer and e-mail address, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict whether a user is a web or a machine learning developer (could also be both or neither).
B.1.2. MUSAE-Facebook
@@
This is a large Facebook page-page network introduced in MUSAE. Nodes represent verified pages on Facebook, and edges are mutual likes. HyperAug constructs the virtual hyperedges as mutually liked page groups that contain at least 3 pages (i.e., maximal cliques with sizes of at least 3). We enable the option to use either the 4,714-dimensional raw node features extracted from the site descriptions, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict which of the four categories a page belongs to: politicians, governmental organizations, television shows and companies.
B.1.3. MUSAE-Twitch
@@
These are six small Twitch user-user networks introduced in MUSAE. Nodes represent gamers on Twitch, and edges are mutual follower relationships between them. HyperAug constructs the virtual hyperedges as mutually following user groups that contain at least 3 gamers (i.e., maximal cliques with sizes of at least 3). We enable the option to use either the 3,170-dimensional raw node features extracted based on the games played and liked, location and streaming habits, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict whether a user streams mature content.
B.1.4. MUSAE-Wiki
@@
These are three Wikipedia page-page networks dataset introduced in MUSAE. Nodes represent articles, and edges represent mutual hyperlinks between them. HyperAug constructs the virtual hyperedges as mutually linked page groups that contain at least 3 pages (i.e., maximal cliques with sizes of at least 3). We enable each dataset to have an option to use either the raw node features extracted based on informative nouns appeared in the text of the Wikipedia articles, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict the average monthly traffic of the web page.
B.2. GRAND Datasets
B.2.1. GRAND-Tissues
@@
We select and build six gene regulatory networks in different tissues (artery aorta, artery coronary, breast, brain, lung, and stomach) from GRAND (Ben Guebila et al., 2022), a public database for gene regulation. Nodes represent gene regulatory elements (Maston et al., 2006) with three distinct types: protein-encoding gene, lncRNA gene (Long et al., 2017), and other regulatory elements. Edges are regulatory effects between genes. We train a CNN (Eraslan et al., 2019) and use it to take the gene sequence as input and create a 4,651-dimensional embedding for each node. HyperAug constructs the virtual hyperedges by grouping nearby genomic elements on the chromosomes, i.e., the genomic elements within 200k base pair distance are grouped as virtual hyperedges. The task is a multi-class classification of gene regulatory elements.
B.2.2. GRAND-Diseases
@@
We select and build four gene regulatory networks in different genetic diseases (leukemia, lung cancer, stomach cancer, and kidney cancer) from GRAND, a public database for gene regulation. Nodes represent gene regulatory elements with three distinct types: protein-encoding gene, lncRNA gene, and other regulatory elements. Edges are regulatory effects between genes. We train a CNN and use it to take the gene sequence as input and create a 4,651-dimensional embedding for each node. HyperAug constructs the virtual hyperedges by grouping nearby genomic elements on the chromosomes, i.e., the genomic elements within 200k base pair distance are grouped as virtual hyperedges. The task is a multi-class classification of gene regulatory elements.
B.3. Amazon Datasets
@@
The Amazon Computers and Photos datasets are two e-commerce graphs based on the Amazon Product Reviews dataset (McAuley et al., 2015; He and McAuley, 2016; Ni et al., 2019). Nodes represent products, and an edge between two products is established if a user buys or writes reviews for both products. Node features are extracted based on the textual description of the products. HyperAug constructs the virtual hyperedges through the following process: first, it feeds the raw images of the products into a CLIP (Radford et al., 2021) classifier, obtaining a 512-dimensional feature embedding for each image to assist the clustering; it then groups the products whose image embeddings’ pairwise distances are within a certain threshold into the same virtual hyperedges. The task is to predict the sub-category of a product in a multi-class classification setup.
Appendix C Experimental Setup
C.1. Training Details
We run all the experiments on NVIDIA A100 PCIe GPUs with 40GB memory (Sulis), NVIDIA V100 NVLink GPUs with 32GB memory (JADE), and NVIDIA V100 SXM2 GPUs with 16GB memory (Cirrus). Adam (Kingma and Ba, 2015) is used as the optimiser, and CosineAnnealingLR (Gotmare et al., 2019) is used as the learning rate scheduler for all training. For each experiment, the nodes of the graph dataset are randomly split into training, validation, and test sets with a split ratio of 6:2:2. All models are trained for 500 epochs. For node classification tasks, the negative log likelihood loss (NLLLoss) is used as the loss function. For node regression tasks, the mean square error (MSELoss) is used as the loss function. Training typically takes less than 5 minutes for each experiment, when ED-HNN is not incorporated in the model. Due to the significantly larger architecture of ED-HNN, experiments involving it can take up to 2 hours.
C.2. Hyperparameter Settings
We perform a hyperparameter search for the learning rate and dropout rate, while keeping the hidden dimension of the layers fixed. To ensure fair comparison, all evaluated GNNs (GCN, GAT, GraphSAGE, HyperConv, and ED-HNN) share the same hyperparameter combinations. After hyperparameter searching, we adopt the following hyperparameter selections: learning rate = 0.001, hidden dimension = 64, and dropout rate = 0.5. For the additional hyperparameters of ED-HNN, we closely adhere to the hyperparameter settings specified in the original ED-HNN paper (Wang et al., 2023), and set the number of all inner multi-layer perceptrons (MLPs) within ED-HNN to 2.
Appendix D Results
@@
  • Simple-GNNs with HyperAug, using simple-GNNs as feature extraction functions:
    • GCN+HyperAug(GAT)
    • GCN+HyperAug(GraphSAGE)
    • GAT+HyperAug(GraphSAGE)
@@
    These are used as baselines of ablation studies, as described in Section 4.2.2;
  • A hyper-GNN with HyperAug, using a hyper-GNN as the feature extraction function:
    • HyperConv+HyperAug(ED-HNN)
@@
    This is intended to assess whether the auxiliary features produced by HyperAug can be used to augment themselves, and improve the node prediction performance using only hyperedge information;
@@
  • Simple-GNNs with HyperAug, using hyper-GNNs as feature extraction functions:
    • GCN+HyperAug(HyperConv)
    • GCN+HyperAug(ED-HNN)
    • GAT+HyperAug(HyperConv)
    • GAT+HyperAug(ED-HNN)
    • GraphSAGE+HyperAug(GraphSAGE)
    • GraphSAGE+HyperAug(ED-HNN)
    These are the standard usage of HyperAug.
@@
Based on the time and compute resource constraints, each experiment is repeated two to five times with different random seeds. The results are summarised in Tables 8, 9, 11, 12, 13, 14 and 10, organised by the dataset subgroups. The results clearly demonstrate that simple-GNNs with HyperAug using hyper-GNNs as feature extraction functions consistently outperform the vanilla GNN baselines and other HyperAug combinations, underscoring the superior efficacy of incorporating information from higher-order node relations into graph augmentation methods under the HyperAug framework. Among all the HyperAug methods, GraphSAGE+HyperAug(HyperConv) generally stands out as the most effective combination.
Figure 4 summarises the performance of HyperConv (accuracy for prediction tasks, and MSE for regression tasks) with the three GraphSAINT samplers on 21 virtual hypergraphs produced by HyperAug, other than MUSAE-GitHub and MUSAE-Facebook (which are illustrated in Figure 3). HyperAug with the GraphSAINT sampling techniques generally enhance HyperConv’s accuracy across most graphs, and are especially significant for regression tasks like MUSAE-Chameleon, MUSAE-Crocodile, and MUSAE-Squirrel.
Appendix E Licence
@@
The raw data for the MUSAE datasets are licenced under the the GNU General Public Licence, version 3 (GPLv3) [^1] https://www.gnu.org/licenses/gpl-3.0.html. The raw data for the GRAND datasets are licenced under the Creative Commons Attribution-ShareAlike 4.0 International Public Licence (CC BY-SA 4.0) [^2] https://creativecommons.org/licenses/by-sa/4.0/. The raw data for the Amazon datasets are licenced under the Amazon Service licence [^3] https://s3.amazonaws.com/amazon-reviews-pds/LICENSE.txt. Having carefully observed the licence requirements of all data sources and code dependencies, we apply the same licenses to our constructed datasets as those of the raw data respectively.
Appendix F Ethics Statement
@@
All datasets constructed in this project are generated from public open-source datasets, and the original raw data downloaded from the data sources do not contain any personally identifiable information or other sensitive contents. The node prediction tasks for the datasets constructed in this project are designed to ensure that they do not, by any means, lead to discriminations against any social groups. Therefore, we are not aware of any social or ethical concern of HyperAug. Since HyperAug is a general graph augmentation method for representation learning on complex graphs, we also do not foresee any direct application of HyperAug to malicious purposes. However, the community should be aware of any potential negative social and ethical impacts that may arise from their chosen datasets or tasks outside of those provided in this project.
@@
Table 8. Accuracy (%) of the HyperAug variants and vanilla GNN baselines on the MUSAE-Facebook and GitHub datasets.
@@
GCN+HyperAug(GAT) 91.0 $\pm$ 0.1 86.7 $\pm$ 0.1
GCN+HyperAug(GraphSAGE) 93.8 $\pm$ 0.3 87.0 $\pm$ 0.4
GAT+HyperAug(GraphSAGE) 93.7 $\pm$ 0.4 86.8 $\pm$ 0.6
HyperConv+HyperAug(ED-HNN) 94.2 $\pm$ 0.1 86.9 $\pm$ 0.3
GCN+HyperAug(HyperConv) 89.8 $\pm$ 0.0 87.2 $\pm$ 0.0
GCN+HyperAug(ED-HNN) 92.6 $\pm$ 0.7 87.4 $\pm$ 0.3
GAT+HyperAug(HyperConv) 90.5 $\pm$ 0.0 86.0 $\pm$ 0.2
GAT+HyperAug(ED-HNN) 93.4 $\pm$ 0.1 87.3 $\pm$ 0.1
GraphSAGE+HyperAug(HyperConv) 93.6 $\pm$ 0.0 87.3 $\pm$ 0.2
GraphSAGE+HyperAug(ED-HNN) 93.2 $\pm$ 0.7 87.4 $\pm$ 0.1
Table 9. Accuracy (%) of the HyperAug variants and vanilla GNN baselines on the MUSAE-Twitch datasets.
@@
GCN+HyperAug(GAT) 72.7 $\pm$ 0.1 62.3 $\pm$ 0.1 66.2 $\pm$ 0.1 61.2 $\pm$ 0.2 68.6 $\pm$ 0.8 74.5 $\pm$ 0.1
GCN+HyperAug(GraphSAGE) 71.8 $\pm$ 2.5 61.2 $\pm$ 0.0 68.2 $\pm$ 1.6 59.2 $\pm$ 2.1 70.0 $\pm$ 1.9 74.8 $\pm$ 0.0
GAT+HyperAug(GraphSAGE) 70.8 $\pm$ 0.7 61.4 $\pm$ 0.0 67.7 $\pm$ 0.5 59.3 $\pm$ 0.2 68.0 $\pm$ 2.1 74.9 $\pm$ 0.4
HyperConv+HyperAug(ED-HNN) 71.8 $\pm$ 1.0 61.2 $\pm$ 0.3 68.3 $\pm$ 0.3 59.3 $\pm$ 0.7 67.9 $\pm$ 0.2 75.1 $\pm$ 0.3
GCN+HyperAug(HyperConv) 72.9 $\pm$ 0.1 62.6 $\pm$ 0.0 65.7 $\pm$ 0.1 60.7 $\pm$ 0.1 69.6 $\pm$ 0.4 74.4 $\pm$ 0.0
GCN+HyperAug(ED-HNN) 72.8 $\pm$ 1.4 63.4 $\pm$ 2.1 67.9 $\pm$ 0.9 62.2 $\pm$ 1.0 70.6 $\pm$ 3.7 75.2 $\pm$ 0.4
GAT+HyperAug(HyperConv) 71.4 $\pm$ 0.1 62.2 $\pm$ 0.0 65.4 $\pm$ 0.1 60.8 $\pm$ 0.3 67.2 $\pm$ 0.3 74.4 $\pm$ 0.0
GAT+HyperAug(ED-HNN) 72.1 $\pm$ 2.3 63.2 $\pm$ 3.1 68.5 $\pm$ 0.2 61.0 $\pm$ 0.6 68.7 $\pm$ 2.8 75.3 $\pm$ 0.2
GraphSAGE+HyperAug(HyperConv) 71.8 $\pm$ 1.0 63.3 $\pm$ 2.2 68.3 $\pm$ 0.3 61.1 $\pm$ 0.2 69.9 $\pm$ 2.4 75.1 $\pm$ 0.3
GraphSAGE+HyperAug(ED-HNN) 70.5 $\pm$ 0.8 63.3 $\pm$ 2.4 68.1 $\pm$ 0.8 60.7 $\pm$ 0.0 68.9 $\pm$ 0.2 75.2 $\pm$ 0.3
Table 10. MSE ($\downarrow$) of the HyperAug variants and vanilla GNN baselines on the MUSAE-Wiki datasets.
Method Squirrel Crocodile Chameleon
GCN 7.319 $\pm$ 0.000 8.761 $\pm$ 0.001 6.779 $\pm$ 0.005
GAT 6.557 $\pm$ 0.154 8.093 $\pm$ 0.054 6.249 $\pm$ 0.261
HyperConv 7.230 $\pm$ 0.002 8.706 $\pm$ 0.000 6.712 $\pm$ 0.001
GCN+HyperAug(GAT) 7.554 $\pm$ 0.029 4.827 $\pm$ 0.087 5.203 $\pm$ 0.147
GCN+HyperAug(HyperConv) 7.313 $\pm$ 0.007 4.851 $\pm$ 0.014 5.515 $\pm$ 0.008
GAT+HyperAug(HyperConv) 6.049 $\pm$ 0.204 4.875 $\pm$ 0.031 4.665 $\pm$ 0.050
Table 11. Accuracy (%) of the HyperAug variants and vanilla GNN baselines on the GRAND-Tissues datasets.
@@
GCN+HyperAug(GAT) 62.7 $\pm$ 0.0 64.1 $\pm$ 0.0 62.6 $\pm$ 0.1 62.5 $\pm$ 0.3 62.7 $\pm$ 0.1 62.6 $\pm$ 0.1
GCN+HyperAug(GraphSAGE) 64.6 $\pm$ 0.3 65.9 $\pm$ 2.2 65.3 $\pm$ 0.3 65.3 $\pm$ 0.9 64.0 $\pm$ 0.1 63.8 $\pm$ 1.6
GAT+HyperAug(GraphSAGE) 66.4 $\pm$ 0.1 66.3 $\pm$ 1.8 65.4 $\pm$ 0.2 65.1 $\pm$ 0.8 65.0 $\pm$ 0.1 65.0 $\pm$ 2.2
HyperConv+HyperAug(ED-HNN) 65.9 $\pm$ 0.0 66.2 $\pm$ 1.1 66.7 $\pm$ 1.0 66.1 $\pm$ 1.6 67.0 $\pm$ 0.7 66.8 $\pm$ 2.7
GCN+HyperAug(HyperConv) 64.7 $\pm$ 0.3 66.0 $\pm$ 0.3 65.2 $\pm$ 0.6 63.7 $\pm$ 0.2 65.4 $\pm$ 0.4 65.4 $\pm$ 0.1
GCN+HyperAug(ED-HNN) 64.3 $\pm$ 0.2 66.0 $\pm$ 2.0 66.1 $\pm$ 1.0 66.7 $\pm$ 2.6 67.9 $\pm$ 1.3 66.9 $\pm$ 0.0
GAT+HyperAug(HyperConv) 64.9 $\pm$ 0.1 66.4 $\pm$ 0.1 65.7 $\pm$ 0.1 64.5 $\pm$ 0.1 66.2 $\pm$ 0.2 65.6 $\pm$ 0.2
GAT+HyperAug(ED-HNN) 65.3 $\pm$ 2.0 66.3 $\pm$ 1.1 65.6 $\pm$ 1.7 65.5 $\pm$ 2.4 68.3 $\pm$ 0.7 66.3 $\pm$ 0.8
GraphSAGE+HyperAug(HyperConv) 66.9 $\pm$ 0.9 67.9 $\pm$ 0.0 67.5 $\pm$ 0.6 66.6 $\pm$ 1.6 67.4 $\pm$ 0.7 68.1 $\pm$ 0.3
GraphSAGE+HyperAug(ED-HNN) 66.2 $\pm$ 0.5 67.1 $\pm$ 2.5 64.7 $\pm$ 0.8 66.0 $\pm$ 0.2 67.9 $\pm$ 0.5 66.9 $\pm$ 0.5
Table 12. Accuracy (%) of the HyperAug variants and vanilla GNN baselines on the GRAND-Diseases datasets.

### v2

@@ Appendix A Full Statistics of Evaluation Datasets
Table 7: Statistics of all 26 datasets built for TopoAug evaluation.
@@ Appendix B Dataset and Task Specifications
B.1 MUSAE Datasets
B.1.1 MUSAE-GitHub
@@
This is a large social network of GitHub developers introduced in MUSAE [29]. Nodes represent developers on GitHub, and edges are mutual follower relationships. TopoAug constructs the virtual hyperedges as mutually following developer groups that contain at least 3 developers (i.e., maximal cliques with sizes of at least 3). We enable the option to use either the 4,005-dimensional raw node features extracted based on the location, repositories starred, employer and e-mail address, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict whether a user is a web or a machine learning developer (could also be both or neither).
B.1.2 MUSAE-Facebook
@@
This is a large Facebook page-page network introduced in MUSAE. Nodes represent verified pages on Facebook, and edges are mutual likes. TopoAug constructs the virtual hyperedges as mutually liked page groups that contain at least 3 pages (i.e., maximal cliques with sizes of at least 3). We enable the option to use either the 4,714-dimensional raw node features extracted from the site descriptions, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict the category a page belongs to: politicians, governmental organizations, television shows and companies.
B.1.3 MUSAE-Twitch
@@
These are six small Twitch user-user networks introduced in MUSAE. Nodes represent gamers on Twitch, and edges are mutual follower relationships between them. TopoAug constructs the virtual hyperedges as mutually following user groups that contain at least 3 gamers (i.e., maximal cliques with sizes of at least 3). We enable the option to use either the 3,170-dimensional raw node features extracted based on the games played and liked, location and streaming habits, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict whether a user streams mature content.
B.1.4 MUSAE-Wiki
@@
These are three Wikipedia page-page networks dataset introduced in MUSAE. Nodes represent articles, and edges represent mutual hyperlinks between them. TopoAug constructs the virtual hyperedges as mutually linked page groups that contain at least 3 pages (i.e., maximal cliques with sizes of at least 3). We enable each dataset to have an option to use either the raw node features extracted based on informative nouns appeared in the text of the Wikipedia articles, or the 128-dimensional preprocessed node embeddings by MUSAE. The task is to predict the average monthly traffic of the web page.
B.2 GRAND Datasets
B.2.1 GRAND-Tissues
@@
We select and build six gene regulatory networks in different tissues (artery aorta, artery coronary, breast, brain, lung, and stomach) from GRAND [2], a public database for gene regulation. Nodes represent gene regulatory elements [23] with three distinct types: protein-encoding gene, lncRNA gene [22], and other regulatory elements. Edges are regulatory effects between genes. We train a CNN [8] and use it to take the gene sequence as input and create a 4,651-dimensional embedding for each node. TopoAug constructs the virtual hyperedges by grouping nearby genomic elements on the chromosomes, i.e., the genomic elements within 200k base pair distance are grouped as virtual hyperedges. The task is a multi-class classification of gene regulatory elements.
B.2.2 GRAND-Diseases
@@
We select and build four gene regulatory networks in different genetic diseases (leukemia, lung cancer, stomach cancer, and kidney cancer) from GRAND, a public database for gene regulation. Nodes represent gene regulatory elements with three distinct types: protein-encoding gene, lncRNA gene, and other regulatory elements. Edges are regulatory effects between genes. We train a CNN and use it to take the gene sequence as input and create a 4,651-dimensional embedding for each node. TopoAug constructs the virtual hyperedges by grouping nearby genomic elements on the chromosomes, i.e., the genomic elements within 200k base pair distance are grouped as virtual hyperedges. The task is a multi-class classification of gene regulatory elements.
B.3 Amazon Datasets
@@
The Amazon Computers and Photos datasets are two e-commerce graphs based on the Amazon Product Reviews dataset [24, 16, 26]. Nodes represent products, and an edge between two products is established if a user buys or writes reviews for both products. Node features are extracted based on the textual description of the products. TopoAug constructs the virtual hyperedges through the following process: first, it feeds the raw images of the products into a CLIP [27] classifier, obtaining a 512-dimensional feature embedding for each image to assist the clustering; it then groups the products whose image embeddings’ pairwise distances are within a certain threshold into the same virtual hyperedges. The task is to predict the sub-category of a product in a multi-class classification setup.
Appendix C Results
@@
  • Simple-GNNs with TopoAug, using simple-GNNs as auxiliary functions (these are used as baselines of ablation studies, as described in Section 4.3.2):
    • GCN+TopoAug(GAT)
    • GCN+TopoAug(GraphSAGE)
    • GAT+TopoAug(GraphSAGE)
  • A hyper-GNN with TopoAug, using another hyper-GNN as the auxiliary function:
    • HyperConv+TopoAug(ED-HNN)
@@
    This is intended to assess whether the auxiliary features produced by TopoAug can be used to augment themselves, and improve the node prediction performance using only hyperedge information;
@@
  • Simple-GNNs with TopoAug, using hyper-GNNs as auxiliary functions (these are the standard usage of TopoAug):
    • GCN+TopoAug(HyperConv)
    • GCN+TopoAug(ED-HNN)
    • GAT+TopoAug(HyperConv)
    • GAT+TopoAug(ED-HNN)
    • GraphSAGE+TopoAug(GraphSAGE)
    • GraphSAGE+TopoAug(ED-HNN)
@@
Based on the time and compute resource constraints, each experiment is repeated two to five times with different random seeds. The results are summarised in Tables 8, 9, 10, 11, 12, 13 and 6, organised by the dataset subgroups. The results clearly demonstrate that simple-GNNs with TopoAug using hyper-GNNs as auxiliary functions consistently outperform the vanilla GNN baselines and other TopoAug combinations, underscoring the superior efficacy of incorporating information from higher-order node relations into graph augmentation methods under the TopoAug framework. Among all the TopoAug methods, GraphSAGE+TopoAug(HyperConv) generally stands out as the most effective combination.
Appendix D Licence
@@
The raw data for the MUSAE datasets are licenced under the the GNU General Public Licence, version 3 (GPLv3)11footnotemark: 1. The raw data for the GRAND datasets are licenced under the Creative Commons Attribution-ShareAlike 4.0 International Public Licence (CC BY-SA 4.0) [^2] https://creativecommons.org/licenses/by-sa/4.0/. The raw data for the Amazon datasets are licenced under the Amazon Service licence [^3] https://s3.amazonaws.com/amazon-reviews-pds/LICENSE.txt. Having carefully observed the licence requirements of all data sources and code dependencies, we apply the same licenses to our constructed datasets as those of the raw data respectively.
Appendix E Ethics Statement
@@
All datasets constructed in this project are generated from public open-source datasets, and the original raw data downloaded from the data sources do not contain any personally identifiable information or other sensitive contents. The node prediction tasks for the datasets constructed in this project are designed to ensure that they do not, by any means, lead to discriminations against any social groups. Therefore, we are not aware of any social or ethical concern of TopoAug. Since TopoAug is a general graph augmentation method for representation learning on complex graphs, we also do not foresee any direct application of TopoAug to malicious purposes. However, the community should be aware of any potential negative social and ethical impacts that may arise from their chosen datasets or tasks outside of those provided in this project.
@@
Table 8: Accuracy (%) of the TopoAug variants and vanilla GNN baselines on the MUSAE-Facebook and GitHub datasets.
@@
GCN+TopoAug(GAT) 91.0 $\pm$ 0.1 86.7 $\pm$ 0.1
GCN+TopoAug(GraphSAGE) 93.8 $\pm$ 0.3 87.0 $\pm$ 0.4
GAT+TopoAug(GraphSAGE) 93.7 $\pm$ 0.4 86.8 $\pm$ 0.6
HyperConv+TopoAug(ED-HNN) 94.2 $\pm$ 0.1 86.9 $\pm$ 0.3
GCN+TopoAug(HyperConv) 89.8 $\pm$ 0.0 87.2 $\pm$ 0.0
GCN+TopoAug(ED-HNN) 92.6 $\pm$ 0.7 87.4 $\pm$ 0.3
GAT+TopoAug(HyperConv) 90.5 $\pm$ 0.0 86.0 $\pm$ 0.2
GAT+TopoAug(ED-HNN) 93.4 $\pm$ 0.1 87.3 $\pm$ 0.1
GraphSAGE+TopoAug(HyperConv) 93.6 $\pm$ 0.0 87.3 $\pm$ 0.2
GraphSAGE+TopoAug(ED-HNN) 93.2 $\pm$ 0.7 87.4 $\pm$ 0.1
Table 9: Accuracy (%) of the TopoAug variants and vanilla GNN baselines on the MUSAE-Twitch datasets.
@@
GCN+TopoAug(GAT) 72.7 $\pm$ 0.1 62.3 $\pm$ 0.1 66.2 $\pm$ 0.1 61.2 $\pm$ 0.2 68.6 $\pm$ 0.8 74.5 $\pm$ 0.1
GCN+TopoAug(GraphSAGE) 71.8 $\pm$ 2.5 61.2 $\pm$ 0.0 68.2 $\pm$ 1.6 59.2 $\pm$ 2.1 70.0 $\pm$ 1.9 74.8 $\pm$ 0.0
GAT+TopoAug(GraphSAGE) 70.8 $\pm$ 0.7 61.4 $\pm$ 0.0 67.7 $\pm$ 0.5 59.3 $\pm$ 0.2 68.0 $\pm$ 2.1 74.9 $\pm$ 0.4
HyperConv+TopoAug(ED-HNN) 71.8 $\pm$ 1.0 61.2 $\pm$ 0.3 68.3 $\pm$ 0.3 59.3 $\pm$ 0.7 67.9 $\pm$ 0.2 75.1 $\pm$ 0.3
GCN+TopoAug(HyperConv) 72.9 $\pm$ 0.1 62.6 $\pm$ 0.0 65.7 $\pm$ 0.1 60.7 $\pm$ 0.1 69.6 $\pm$ 0.4 74.4 $\pm$ 0.0
GCN+TopoAug(ED-HNN) 72.8 $\pm$ 1.4 63.4 $\pm$ 2.1 67.9 $\pm$ 0.9 62.2 $\pm$ 1.0 70.6 $\pm$ 3.7 75.2 $\pm$ 0.4
GAT+TopoAug(HyperConv) 71.4 $\pm$ 0.1 62.2 $\pm$ 0.0 65.4 $\pm$ 0.1 60.8 $\pm$ 0.3 67.2 $\pm$ 0.3 74.4 $\pm$ 0.0
GAT+TopoAug(ED-HNN) 72.1 $\pm$ 2.3 63.2 $\pm$ 3.1 68.5 $\pm$ 0.2 61.0 $\pm$ 0.6 68.7 $\pm$ 2.8 75.3 $\pm$ 0.2
GraphSAGE+TopoAug(HyperConv) 71.8 $\pm$ 1.0 63.3 $\pm$ 2.2 68.3 $\pm$ 0.3 61.1 $\pm$ 0.2 69.9 $\pm$ 2.4 75.1 $\pm$ 0.3
GraphSAGE+TopoAug(ED-HNN) 70.5 $\pm$ 0.8 63.3 $\pm$ 2.4 68.1 $\pm$ 0.8 60.7 $\pm$ 0.0 68.9 $\pm$ 0.2 75.2 $\pm$ 0.3
Table 10: Accuracy (%) of the TopoAug variants and vanilla GNN baselines on the GRAND-Tissues datasets.
@@
GCN+TopoAug(GAT) 62.7 $\pm$ 0.0 64.1 $\pm$ 0.0 62.6 $\pm$ 0.1 62.5 $\pm$ 0.3 62.7 $\pm$ 0.1 62.6 $\pm$ 0.1
GCN+TopoAug(GraphSAGE) 64.6 $\pm$ 0.3 65.9 $\pm$ 2.2 65.3 $\pm$ 0.3 65.3 $\pm$ 0.9 64.0 $\pm$ 0.1 63.8 $\pm$ 1.6
GAT+TopoAug(GraphSAGE) 66.4 $\pm$ 0.1 66.3 $\pm$ 1.8 65.4 $\pm$ 0.2 65.1 $\pm$ 0.8 65.0 $\pm$ 0.1 65.0 $\pm$ 2.2
HyperConv+TopoAug(ED-HNN) 65.9 $\pm$ 0.0 66.2 $\pm$ 1.1 66.7 $\pm$ 1.0 66.1 $\pm$ 1.6 67.0 $\pm$ 0.7 66.8 $\pm$ 2.7
GCN+TopoAug(HyperConv) 64.7 $\pm$ 0.3 66.0 $\pm$ 0.3 65.2 $\pm$ 0.6 63.7 $\pm$ 0.2 65.4 $\pm$ 0.4 65.4 $\pm$ 0.1
GCN+TopoAug(ED-HNN) 64.3 $\pm$ 0.2 66.0 $\pm$ 2.0 66.1 $\pm$ 1.0 66.7 $\pm$ 2.6 67.9 $\pm$ 1.3 66.9 $\pm$ 0.0
GAT+TopoAug(HyperConv) 64.9 $\pm$ 0.1 66.4 $\pm$ 0.1 65.7 $\pm$ 0.1 64.5 $\pm$ 0.1 66.2 $\pm$ 0.2 65.6 $\pm$ 0.2
GAT+TopoAug(ED-HNN) 65.3 $\pm$ 2.0 66.3 $\pm$ 1.1 65.6 $\pm$ 1.7 65.5 $\pm$ 2.4 68.3 $\pm$ 0.7 66.3 $\pm$ 0.8
GraphSAGE+TopoAug(HyperConv) 66.9 $\pm$ 0.9 67.9 $\pm$ 0.0 67.5 $\pm$ 0.6 66.6 $\pm$ 1.6 67.4 $\pm$ 0.7 68.1 $\pm$ 0.3
GraphSAGE+TopoAug(ED-HNN) 66.2 $\pm$ 0.5 67.1 $\pm$ 2.5 64.7 $\pm$ 0.8 66.0 $\pm$ 0.2 67.9 $\pm$ 0.5 66.9 $\pm$ 0.5
Table 11: Accuracy (%) of the TopoAug variants and vanilla GNN baselines on the GRAND-Diseases datasets.